OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 7) and (–1, 3)? (1 point) 2/7 3/4 4/3 1/3

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (baldymcgee6):

Well, we do have the slope formula.. this will probably do the trick :)

OpenStudy (anonymous):

I am not sure how to do it I completely forgot :\

OpenStudy (baldymcgee6):

that's okay, i'll refresh your memory.. When we have two sets of co-ordinates, like they gave us, we can use the formula:\[\LARGE m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (baldymcgee6):

For example: From the points given, (2, 7) and (–1, 3) (x1,y1) and (x2, y2)

OpenStudy (baldymcgee6):

Do ya get it?

OpenStudy (anonymous):

then what do you do next?

OpenStudy (anonymous):

I kinda get it but not much :\

OpenStudy (baldymcgee6):

\[\LARGE m=\frac{ y_2-y_1 }{ x_2-x_1 } = \frac{ 3-7 }{ -1-2 } = \frac{ 4 }{ 3 }\] Simply sub in the points given and do the math :)

OpenStudy (anonymous):

oooooh!

OpenStudy (baldymcgee6):

:D the sound of learning

OpenStudy (baldymcgee6):

@...help... do you understand it?

OpenStudy (anonymous):

Yep I was just looking over it :P THANKS A BUNCH

OpenStudy (baldymcgee6):

No problem!

OpenStudy (anonymous):

:)